# Jinhong Li1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

### v2

@@ Abstract
Zoned Namespace (ZNS) defines a new abstraction for host software to flexibly manage storage in flash-based SSDs as append-only zones. It also provides a Zone Append primitive to further boost the write performance of ZNS SSDs by exploiting intra-zone parallelism. However, making Zone Append effective for reliable and scalable storage, in the form of a RAID array of multiple ZNS SSDs, is non-trivial since Zone Append offloads address management to ZNS SSDs and requires hosts to dedicatedly manage RAID stripes across multiple drives. We propose ZapRAID, a high-performance log-structured RAID system for ZNS SSDs by carefully exploiting Zone Append to achieve high write parallelism and lightweight stripe management. ZapRAID adopts a group-based data layout with a coarse-grained ordering across multiple groups of stripes, such that it can use small-size metadata for stripe management on a per-group basis under Zone Append. It further adopts hybrid data management to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Write and Zone Append primitives. We implement ZapRAID as a user-space block device, and evaluate ZapRAID using microbenchmarks, trace-driven experiments, and real-application experiments. Our evaluation results show that ZapRAID achieves high write throughput and maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
@@ 1 Introduction
Flash-based solid-state drives (SSDs) have been widely deployed as they have better performance and reliability than hard-disk drives (HDDs). Conventional SSDs expose a block interface, implemented inside the flash translation layer (FTL), to a host for accessing flash storage. However, such a block interface also introduces costly internal flash management operations (e.g., device-level garbage collection and address translation) that lead to unpredictable I/O performance [34] and reduced flash endurance [21]. The recently proposed NVMe Zoned Namespace (ZNS) interface [15, 21] abstracts flash-based SSDs as append-only zones, such that the FTL and its costly operations are eliminated and the storage management is now shifted to the host. Compared with conventional SSDs, ZNS SSDs are shown to achieve higher write throughput, lower tail read latencies, and less device-level DRAM usage [21].
ZNS SSDs support two write primitives, namely Zone Write and Zone Append, where Zone Append provides opportunities for further performance gains by exploiting intra-zone parallelism. Specifically, a ZNS SSD tracks per-zone write pointers, such that any write to a zone must specify the same offset indicated by the write pointer. Zone Write requires the host to specify the block address when writing a block as in conventional SSDs. This provides backward compatibility with existing applications. However, in order to match the offset of a Zone Write with the on-device write pointer of a zone, the host can only issue one request to each zone at a time, thereby limiting intra-zone parallelism. In contrast, Zone Append eliminates the need of the host to specify block addresses in writes by fully offloading address management to ZNS SSDs. In Zone Append, the host only specifies the zone to which a write is issued, and the ZNS SSD returns the address to the host upon write completion. Thus, the host can issue multiple Zone Append commands within one zone to exploit the intra-zone parallelism for improved write performance.
@@
We make the source code available on https://github.com/adslabcuhk/zapraid for review.
@@ 2.1 Zoned Namespace (ZNS)
The Zoned Namespace (ZNS) interface [15, 21] is a new NVMe Command Set [4] that exposes flash-based SSD internals to the host for host-level storage management. Figure 1 shows the architecture of a ZNS SSD based on the Western Digital Ultrastar DC ZN540 model, which we use in this paper and also previous studies (e.g., [20, 45, 49, 69, 32]). The ZNS interface abstracts a ZNS SSD as append-only zones, each of which has a maximum size called the zone capacity. Each zone organizes data in logical blocks (or blocks in short in this paper), and each block is mapped to a physical flash page. Each flash page also has an out-of-band area for metadata storage (of size tens of bytes) located alongside the page content. Note that the out-of-band area is also available in general NVMe SSDs, and its space is not included in the total zone capacity. In this paper, we assume that the block size is 4 KiB and the out-of-band area size is 64 bytes, both of which can be configured in the ZN540 model when a ZNS SSD is formatted.
@@ 2.2 Zone Write versus Zone Append
Recall from §1 that ZNS provides the Zone Write and Zone Append primitives for a host to issue writes to a zone. Their key difference is that Zone Write requires the host to issue writes serially and specify the offset in each write, while Zone Append offloads address assignment to ZNS SSDs and enables the host to issue concurrent writes that exploit intra-zone parallelism. Note that ZNS can only have one outstanding Zone Write command issued to a zone at a time, and the host cannot execute new Zone Write or Zone Append commands in the same zone if there exists any outstanding Zone Write command that is being executed; otherwise, the new commands may modify the write pointer that is being updated by the outstanding Zone Write command.
We conduct evaluation on a single ZN540 ZNS SSD [14] in our testbed (see §5.1 for testbed details) to show how Zone Append improves write performance over Zone Write. We issue write requests of 4 KiB, 8 KiB, and 16 KiB to the ZNS SSD. We also vary the number of open zones in the ZNS SSD. We write a total of 512 GiB of data and measure the write throughput. We report the average throughput results over five runs, with the error bars showing the 95% confidence intervals under the Student’s t-distribution (note that some error bars may be invisible due to small derivations). Figure 2 shows the write throughput results.
@@
(a) 4-KiB requests (b) 8-KiB requests (c) 16-KiB requests
@@
We first consider the case when the number of open zones is one. In this case, we issue writes to a single zone at a time; if the zone is full, we issue writes to a different zone. Zone Write achieves a write throughput of 337.6 MiB/s, 613.6 MiB/s, and 1,050.0 MiB/s for 4-KiB, 8-KiB, and 16-KiB requests, respectively. Recall that a ZNS SSD can have only one outstanding Zone Write at any time. We find that the write throughput of Zone Write cannot further increase even if we issue multiple concurrent write requests of Zone Write. In contrast, Zone Append achieves a write throughput of 541.5 MiB/s, 1,026.6 MiB/s, and 1,050.1 MiB/s for 4-KiB, 8-KiB, and 16-KiB requests, respectively with four concurrent write requests of Zone Append [^1] The ZN540 specification [14] states that the random write throughput can reach 2,000 MiB/s. This number corresponds to the best-case performance and is obtained when large requests (128 KiB) are issued to all open zones.. This shows that Zone Append can increase the write throughput via intra-zone parallelism.
@@
Note that in this single-zone experiment, even if we issue more than four concurrent write requests, the write throughput of Zone Append for 4-KiB and 8-KiB requests does not further increase as it already saturates intra-zone parallelism for this specific ZN540 model. Also, Zone Write already achieves the maximum write throughput of a zone in ZN540 for 16-KiB requests, so Zone Append does not see more growth.
Writes to multiple open zones.
@@
We further consider the case when the number of open zones is larger than one, so as to examine how inter-zone parallelism affects the effectiveness of Zone Write and Zone Append. We create multiple open zones and issue writes to the open zones in parallel. We set the number of concurrent write requests for each open zone as one and four for Zone Write and Zone Append, respectively, based on the previous results in the single-zone experiment. From Figure 2, we observe that the write throughput of Zone Write is generally higher when there are more open zones (e.g., its write throughput reaches 777.1 MiB/s under six open zones for 4-KiB write requests (Figure 2(a))). On the other hand, Zone Append can only achieve a maximum write throughput of 577.5 MiB/s under two open zones for 4-KiB requests (Figure 2(a)). A possible reason is that the current firmware implementation of Zone Append is more computationally intensive when writes are issued to more open zones in the ZN540 model, so Zone Append has even lower write throughput than Zone Write under a larger number of open zones due to the limited computational power in existing hardware. We conjecture that such a limitation can be addressed in future ZNS SSD models. When the request size is 8 KiB (Figure 2(b)), Zone Write can scale to 1,430.7 MiB/s under eight open zones, while Zone Append can only achieve a maximum write throughput of 1,058.6 MiB/s under two open zones. When the request size is 16 KiB (Figure 2(c)), Zone Write and Zone Append have very small performance differences (e.g., up to around 1,750 MiB/s under eight open zones), implying large writes can effectively exploit inter-zone parallelism in both Zone Write and Zone Append.
Main observations.
@@
We observe that for small writes (e.g., of size 4 KiB or 8 KiB), Zone Append achieves higher write throughput than Zone Write in the single-zone setting by exploiting intra-zone parallelism, while Zone Write scales better and achieves higher write throughput than Zone Append when the number of open zones increases by better exploiting inter-zone parallelism. For large writes (e.g., of size 16 KiB), both Zone Write and Zone Append have similar write throughput for different numbers of open zones. A RAID system based on ZNS SSDs should adequately combine Zone Append and Zone Write to accommodate the scenarios with different numbers of open zones. Even though Zone Append only shows high performance when write requests are issued to one or a few open zones, it is still beneficial for the following scenarios. First, shared storage systems need performance isolation [25, 35, 41]. Specifically, since a zone is typically mapped to only a subset of physical flash channels or chips [19], a shared storage system based on ZNS SSDs should allocate only a limited number of open zones to applications that do not have high write bandwidth demands. This isolates the write traffic of the applications to a subset of flash channels and limits their impact on other co-located applications. Second, it is common to have mixed request sizes in many real-world workloads. For example, practical cloud block storage workloads have 60% of write requests that are at most 4 KiB and 25% of write requests that are at least 16 KiB [55]. By separating small and large writes into different open zones (§3.3), we can use Zone Append and Zone Write for different groups of writes to improve the overall performance (§5.2). Thus, Zone Append can help such applications achieve high write performance through intra-zone parallelism and sustain the bursts of a large number of write requests if necessary, even with only one or a few open zones per ZNS SSD being used.
@@ 2.3 Log-structured RAID (Log-RAID)
By writing blocks as new stripes in an append-only manner, Log-RAID needs to perform garbage collection to reclaim the space from stale blocks. When garbage collection is triggered (say, when the available space drops below some threshold), Log-RAID selects a sealed segment by some policy (e.g., using a greedy algorithm to select the one with the most fraction of stale blocks), rewrites all non-stale blocks into a new open segment, and releases the space of the selected sealed segment for reuse.
@@ 3.1 Design Overview
Architecture. We first introduce the architecture of ZapRAID and show how it augments Log-RAID on an array of multiple ZNS SSDs, as shown in Figure 3. ZapRAID exposes a block-level volume that supports random reads/writes. We assume that the block size is 4 KiB. A user application can issue reads/writes of an arbitrary number of blocks, each of which is identified by a logical block address (LBA). Each drive corresponds to a ZNS SSD, and ZapRAID maps each contiguous area of a drive under Log-RAID to a zone, which only allows sequential writes. Let $Z$ be the total number of zones in a drive, so there are $Z$ segments in a ZNS SSD array. For example, a 4-TiB ZN540 ZNS SSD [14] configures $Z=$ 3,690 zones. Each of the drives, segments, zones, and stripes is associated with an identifier (ID).
@@
Segment organization.
Each segment corresponds to $k+m$ zones in $k+m$ drives. In ZapRAID, each segment comprises three regions that span across $k+m$ drives: the header region and the footer region for keeping metadata for crash consistency (§3.4), and the data region for storing data and parity chunks. Specifically, the header region stores the RAID scheme (including RAID-5/RAID-6, $k$, and $m$) and the segment-specific information (including the zone IDs of all zones from the $k+m$ drives in the segment, the chunk size, and the stripe group size (defined in §3.2)). The footer region keeps the block metadata, including the LBA, write timestamp, and stripe ID of each block in the data region.
@@
Block metadata.
ZapRAID stores the block metadata for each block in the out-of-band area of the corresponding flash page (§2.1) for persistence. Each block in a data chunk has its LBA, write timestamp, and stripe ID as its block metadata. ZapRAID provides fault tolerance for the block metadata: for LBAs and write timestamps, ZapRAID generates parity-based redundancy for them from all blocks of the data chunks in the same stripe and stores the parity results in the block metadata of the blocks of the parity chunks, while for stripe IDs, ZapRAID replicates them into all the data and parity chunks in the same stripe.
@@
In-memory items.
ZapRAID keeps a number of in-memory in-flight stripes for newly written blocks. Each in-flight stripe is kept in memory until all of its $k$ data chunks and $m$ parity chunks are formed and persisted. To maintain durability, ZapRAID acknowledges the writes of an in-flight stripe only after the whole in-flight stripe is persisted (note that acknowledging the write of each block can lead to data loss if a drive storing an acknowledged block fails but the parity chunks are not yet generated).
@@
Offloading L2P table entries to ZNS SSDs.
The L2P table incurs large memory footprints if it keeps all entries in memory. For example, our current implementation supports a RAID array of four 4-TiB ZN540 ZNS SSDs (i.e., a total capacity of 16 TiB) and uses a 4-byte L2P entry for each 4-KiB block. Thus, the total memory size of the L2P table is 16 GiB. Such large L2P table footprints are common in all Log-RAID designs [27, 30, 37, 44]. Some Log-RAID designs keep part of the L2P table entries on disk to mitigate the memory overhead [27, 30].
ZapRAID by default keeps the whole L2P table in memory, yet it also supports the offloading of part of the L2P table entries to ZNS SSDs. It uses the CLOCK algorithm [31] to evict the non-recently used L2P entries to SSDs. Specifically, it divides the L2P table entries into multiple groups, each of which has 1,024 contiguous entries that can be fit into a 4-KiB block (assuming that each entry is of size four bytes). It maintains an in-memory bitmap, in which each bit tracks if the corresponding entry group has been accessed, and a pointer that references the entry group to be considered. It sets the tracked bit to one if any entry in the corresponding entry group has been accessed. Suppose that the L2P table size reaches the memory size limit. ZapRAID then searches for an entry group to evict, starting from the one referenced by the pointer. If the tracked bit of an entry group is one, ZapRAID resets the tracked bit to zero, moves the pointer to the next entry group, and continues the search; otherwise, if the tracked bit of the entry group is zero, ZapRAID selects the entry group to evict, moves the pointer to the next entry group, and stops the search.
ZapRAID maps the entry group into a 4-KiB mapping block. It writes the mapping block together with other user-written blocks, so as to avoid consuming additional open zone resources. It maintains a separate in-memory mapping table to record the PBA of each on-SSD mapping block. Note that the size of the in-memory mapping table is much smaller as each of its entries can track 1,024 L2P table entries (e.g., in our case, the size of the in-memory mapping table is 16 MiB for a RAID array of four 4-TiB ZN540 ZNS SSDs). It also keeps the LBA of the first entry of the mapping block in the LBA field of the block metadata for crash recovery (§3.4).
It is important to distinguish between the mapping blocks and user-written blocks during the crash recovery of the L2P table (§3.4). Thus, ZapRAID marks the least significant bit of the LBA field in the block metadata as one for a mapping block. It is feasible since the LBA field of a user-written block is currently aligned at multiples of 4 KiB and its 12 least significant bits are all zeros.

### v3

# Jinhong Li1    Yiyang Geng1    Qiuping Wang1    Shujie Han2    Patrick P. C. Lee1

@@ Abstract
Zoned Namespace (ZNS) defines a new abstraction for host software to flexibly manage storage in flash-based SSDs as append-only zones. It also provides a Zone Append primitive to further boost the write performance of ZNS SSDs by exploiting intra-zone parallelism. However, making Zone Append effective for reliable and scalable storage, in the form of a RAID array of multiple ZNS SSDs, is non-trivial, since Zone Append offloads address management to ZNS SSDs and requires hosts to specifically manage RAID stripes across multiple drives. We propose ZapRAID, a high-performance log-structured RAID system for ZNS SSDs by carefully exploiting Zone Append to achieve high write parallelism and lightweight stripe management. ZapRAID adopts a group-based data layout with a coarse-grained ordering across multiple groups of stripes, such that it can use small-size metadata for stripe management on a per-group basis under Zone Append. It further adopts hybrid data management to simultaneously achieve intra-zone and inter-zone parallelism through a careful combination of both Zone Write and Zone Append primitives. We implement ZapRAID as a user-space block device, and evaluate ZapRAID using microbenchmarks, trace-driven experiments, and real-application experiments. Our evaluation results show that ZapRAID achieves high write throughput and maintains high performance in normal reads, degraded reads, crash recovery, and full-drive recovery.
@@ 1 Introduction
Flash-based solid-state drives (SSDs) have been widely deployed as they have better performance and reliability than hard-disk drives (HDDs). Conventional SSDs expose a block interface, implemented inside the flash translation layer (FTL), to a host for accessing flash storage. However, such a block interface also introduces costly internal flash management operations (e.g., device-level garbage collection and address translation) that lead to unpredictable I/O performance [19] and reduced flash endurance [5]. The recently proposed NVMe Zoned Namespace (ZNS) interface [77, 5] abstracts flash-based SSDs as append-only zones, so that the FTL and its costly operations are eliminated, and the storage management is now shifted to the host. Compared with conventional SSDs, ZNS SSDs are shown to achieve higher write throughput, lower tail read latencies, and less device-level DRAM usage [5].
ZNS SSDs support two write primitives, namely Zone Write and Zone Append, where Zone Append provides opportunities for further performance gains by exploiting intra-zone parallelism. Specifically, a ZNS SSD tracks per-zone write pointers, such that any write to a zone must specify the same offset indicated by the write pointer. Zone Write requires the host to specify the block address when writing a block, as in conventional SSDs. This provides backward compatibility with existing applications. However, in order to match the offset of a Zone Write with the on-device write pointer of a zone, the host can only issue one request to each zone at a time, thereby limiting intra-zone parallelism. In contrast, Zone Append eliminates the need for the host to specify block addresses in writes by fully offloading address management to ZNS SSDs. In Zone Append, the host only specifies the zone to which a write is issued, and the ZNS SSD returns the address to the host upon write completion. Thus, the host can issue multiple Zone Append commands within one zone to exploit the intra-zone parallelism for improved write performance.
@@
We now make the source code available on https://github.com/adslabcuhk/zapraid.
@@ 2.1 Zoned Namespace (ZNS)
The Zoned Namespace (ZNS) interface [77, 5] is a new NVMe Command Set [50] that exposes flash-based SSD internals to the host for host-level storage management. Figure 1 shows the architecture of a ZNS SSD based on the Western Digital Ultrastar DC ZN540 model, which we use in this paper and also in previous studies (e.g., [4, 31, 35, 60, 16]). The ZNS interface abstracts a ZNS SSD as append-only zones, each of which has a maximum size called the zone capacity. Each zone organizes data in logical blocks (or blocks in short in this paper), and each block is mapped to a physical flash page. Each flash page also has an out-of-band area for metadata storage (of size tens of bytes) located alongside the page content. Note that the out-of-band area is also available in general NVMe SSDs, and its space is not included in the total zone capacity. In this paper, we assume that the block size is 4 KiB and the out-of-band area size is 64 bytes, both of which can be configured in the ZN540 model when a ZNS SSD is formatted.
@@ 2.2 Zone Write versus Zone Append
Recall from §1 that ZNS provides the Zone Write and Zone Append primitives for a host to issue writes to a zone. Their key difference is that Zone Write requires the host to issue writes serially and specify the offset in each write, while Zone Append offloads address assignment to ZNS SSDs and enables the host to issue concurrent writes that exploit intra-zone parallelism. Existing ZNS devices and host software typically allow only one outstanding Zone Write per zone in order to maintain the write pointer invariant; otherwise, concurrent Zone Write and Zone Append commands may violate the sequential-write requirement.
We conduct evaluation on a single ZN540 ZNS SSD [70] in our testbed (see §5.1 for testbed details) to show how Zone Append improves write performance over Zone Write. We issue write requests of 4 KiB, 8 KiB, and 16 KiB to the ZNS SSD. We also vary the number of open zones in the ZNS SSD. We write a total of 512 GiB of data and measure the write throughput. We report the average throughput results over five runs, with the error bars showing the 95% confidence intervals under the Student’s t-distribution (note that some error bars may be invisible due to small deviations). Figure 2 shows the write throughput results.
@@
(a) 4 KiB requests (b) 8 KiB requests (c) 16 KiB requests
@@
Writes to a single open zone. We first consider the case when the number of open zones is one. In this case, we issue writes to a single zone at a time; if the zone is full, we issue writes to a different zone. Zone Write achieves a write throughput of 337.6 MiB/s, 613.6 MiB/s, and 1,050.0 MiB/s for 4 KiB, 8 KiB, and 16 KiB requests, respectively. Recall that a ZNS SSD can have only one outstanding Zone Write at any time. We find that the write throughput of Zone Write cannot further increase even if we issue multiple concurrent write requests of Zone Write. In contrast, Zone Append achieves a write throughput of 541.5 MiB/s, 1,026.6 MiB/s, and 1,050.1 MiB/s for 4 KiB, 8 KiB, and 16 KiB requests, respectively, with four concurrent write requests of Zone Append [^1] The ZN540 specification [70] states that the write throughput can reach 2,000 MiB/s. This number corresponds to the best-case performance and is obtained when large requests (128 KiB) are issued to all open zones.. This shows that Zone Append can increase the write throughput via intra-zone parallelism.
@@
Note that in this single-zone experiment, even if we issue more than four concurrent write requests, the write throughput of Zone Append for 4 KiB and 8 KiB requests does not further increase, as it already saturates intra-zone parallelism for this specific ZN540 model. Also, Zone Write already achieves the maximum write throughput of a zone in ZN540 for 16 KiB requests, so Zone Append does not see more growth.
@@
Writes to multiple open zones. We further consider the case when the number of open zones is larger than one, so as to examine how inter-zone parallelism affects the effectiveness of Zone Write and Zone Append. We create multiple open zones and issue writes to the open zones in parallel. We set the number of concurrent write requests for each open zone as one and four for Zone Write and Zone Append, respectively, based on the previous results in the single-zone experiment. From Figure 2, we observe that the write throughput of Zone Write is generally higher when there are more open zones (e.g., its write throughput reaches 777.1 MiB/s under six open zones for 4 KiB write requests (Figure 2(a))). On the other hand, Zone Append can only achieve a maximum write throughput of 577.5 MiB/s under two open zones for 4 KiB requests (Figure 2(a)). A possible reason is that the current firmware implementation of Zone Append is more computationally intensive when writes are issued to more open zones in the ZN540 model, so Zone Append has even lower write throughput than Zone Write under a larger number of open zones due to the limited computational power in existing hardware. We conjecture that this limitation could be addressed in future ZNS SSD models. When the request size is 8 KiB (Figure 2(b)), Zone Write can scale to 1,430.7 MiB/s under eight open zones, while Zone Append can only achieve a maximum write throughput of 1,058.6 MiB/s under two open zones. When the request size is 16 KiB (Figure 2(c)), Zone Write and Zone Append have very small performance differences (e.g., up to around 1,750 MiB/s under eight open zones), implying large writes can effectively exploit inter-zone parallelism in both Zone Write and Zone Append.
@@
Main observations. We observe that for small writes (e.g., of size 4 KiB or 8 KiB), Zone Append achieves higher write throughput than Zone Write in the single-zone setting by exploiting intra-zone parallelism, while Zone Write scales better and achieves higher write throughput than Zone Append when the number of open zones increases by better exploiting inter-zone parallelism. For large writes (e.g., of size 16 KiB), both Zone Write and Zone Append have similar write throughput for different numbers of open zones. A RAID system based on ZNS SSDs should adequately combine Zone Append and Zone Write to accommodate the scenarios with different numbers of open zones. Even though Zone Append only shows high performance when write requests are issued to one or a few open zones, it is still beneficial for the following scenarios. First, shared storage systems need performance isolation [9, 20, 27]. Specifically, since a zone is typically mapped to only a subset of physical flash channels or chips [3], a shared storage system based on ZNS SSDs should allocate only a limited number of open zones to applications that do not have high write bandwidth demands. This isolates the write traffic of the applications to a subset of flash channels and limits their impact on other co-located applications. Second, it is common to have mixed request sizes in many real-world workloads. For example, practical cloud block storage workloads have 60% of write requests that are at most 4 KiB and 25% of write requests that are at least 16 KiB [41]. By separating small and large writes into different open zones (§3.3), we can use Zone Append and Zone Write for different groups of writes to improve the overall performance (§5.4). Thus, Zone Append can help such applications achieve high write performance through intra-zone parallelism and sustain the bursts of a large number of write requests if necessary, even with only one or a few open zones per ZNS SSD being used.
@@ 2.3 Log-structured RAID (Log-RAID)
By writing blocks as new stripes in an append-only manner, Log-RAID needs to perform garbage collection to reclaim the space from stale blocks. When garbage collection is triggered (say, when the available space drops below some threshold), Log-RAID selects a sealed segment by some policy (e.g., using a greedy algorithm to select the one with the largest fraction of stale blocks), rewrites all non-stale blocks into a new open segment, and releases the space of the selected sealed segment for reuse.
@@ 3.1 Design Overview
Architecture. We first introduce the architecture of ZapRAID and show how it augments Log-RAID on an array of multiple ZNS SSDs, as shown in Figure 3. ZapRAID exposes a block-level volume that supports random reads/writes. We assume that the block size is 4 KiB. A user application can issue reads/writes of an arbitrary number of blocks, each of which is identified by a logical block address (LBA). Each drive corresponds to a ZNS SSD, and ZapRAID maps each contiguous area of a drive under Log-RAID to a zone, which only allows sequential writes. Let $Z$ be the total number of zones in a drive, so there are $Z$ segments in a ZNS SSD array. For example, a 4 TiB ZN540 ZNS SSD [70] configures $Z=$ 3,690 zones. Each of the drives, segments, zones, and stripes is associated with an identifier (ID).
@@
Segment organization. Each segment corresponds to $k+m$ zones in $k+m$ drives. In ZapRAID, each segment comprises three regions that span across $k+m$ drives: the header region and the footer region for keeping metadata for crash consistency (§3.4), and the data region for storing data and parity chunks. Specifically, the header region stores the RAID scheme (including RAID-5/RAID-6, $k$, and $m$) and the segment-specific information (including the zone IDs of all zones from the $k+m$ drives in the segment, the chunk size, and the stripe group size (defined in §3.2)). The footer region keeps the block metadata, including the LBA, write timestamp, and stripe ID of each block in the data region.
@@
Block metadata. ZapRAID stores the block metadata for each block in the out-of-band area of the corresponding flash page (§2.1) for persistence. Each block in a data chunk has its LBA, write timestamp, and stripe ID as its block metadata. ZapRAID provides fault tolerance for the block metadata: for LBAs and write timestamps, ZapRAID generates parity-based redundancy for them from all blocks of the data chunks in the same stripe and stores the parity results in the block metadata of the blocks of the parity chunks, while for stripe IDs, ZapRAID replicates them into all the data and parity chunks in the same stripe.
@@
In-memory items. ZapRAID keeps a number of in-memory in-flight stripes for newly written blocks. Each in-flight stripe is kept in memory until all of its $k$ data chunks and $m$ parity chunks are formed and persisted. To maintain durability, ZapRAID acknowledges the writes of an in-flight stripe only after the whole in-flight stripe is persisted (note that acknowledging the write of each block can lead to data loss if a drive storing an acknowledged block fails, but the parity chunks are not yet generated).
@@
Offloading L2P table entries to ZNS SSDs. The L2P table incurs large memory footprints if it keeps all entries in memory. For example, our current implementation supports a RAID array of four 4 TiB ZN540 ZNS SSDs (i.e., a total capacity of 16 TiB) and uses a 4-byte L2P entry for each 4 KiB block. Thus, the total memory size of the L2P table is 16 GiB. Such large L2P table footprints are common in all Log-RAID designs [11, 14, 23, 30]. Some Log-RAID designs keep part of the L2P table entries on disk to mitigate the memory overhead [11, 14].
ZapRAID by default keeps the whole L2P table in memory, yet it also supports the offloading of part of the L2P table entries to ZNS SSDs. It uses the CLOCK algorithm [15] to evict the non-recently used L2P entries to SSDs. Specifically, it divides the L2P table entries into multiple groups, each of which has 1,024 contiguous entries that can be fit into a 4 KiB block (assuming that each entry is of size four bytes). It maintains an in-memory bitmap, in which each bit tracks if the corresponding entry group has been accessed, and a pointer that references the entry group to be considered. It sets the tracked bit to one if any entry in the corresponding entry group has been accessed. Suppose that the L2P table size reaches the memory size limit. ZapRAID then searches for an entry group to evict, starting from the one referenced by the pointer. If the tracked bit of an entry group is one, ZapRAID resets the tracked bit to zero, moves the pointer to the next entry group, and continues the search; otherwise, if the tracked bit of the entry group is zero, ZapRAID selects the entry group to evict, moves the pointer to the next entry group, and stops the search.
ZapRAID maps the entry group into a 4 KiB mapping block. It writes the mapping block together with other user-written blocks, so as to avoid consuming additional open zone resources. It maintains a separate in-memory mapping table to record the PBA of each on-SSD mapping block. Note that the size of the in-memory mapping table is much smaller, as each of its entries can track 1,024 L2P table entries (e.g., in our case, the size of the in-memory mapping table is 16 MiB for a RAID array of four 4 TiB ZN540 ZNS SSDs). It also keeps the LBA of the first entry of the mapping block in the LBA field of the block metadata for crash recovery (§3.4).
It is important to distinguish between the mapping blocks and user-written blocks during the crash recovery of the L2P table (§3.4). Thus, ZapRAID marks the least significant bit of the LBA field in the block metadata as one for a mapping block. It is feasible since the LBA field of a user-written block is currently aligned at multiples of 4 KiB, and its 12 least significant bits are all zeros.